OpenStudy (anonymous):

i cant find the (SIMPLIFIED) form of the difference quotient for f(x)=sqrt(7x+21)?????

OpenStudy (anonymous):

so far i've got \[\[(\sqrt{7(x+h)+21}-\sqrt{7x+21})\div h \]

OpenStudy (anonymous):

\[\sqrt{7}\]